# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v2

@@ Abstract.
Efficient management of GPU memory is essential for high throughput LLM inference. Prior systems used to reserve KV-cache memory ahead-of-time that resulted in wasted capacity due to internal fragmentation. Inspired by demand paging, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache. This approach eliminates fragmentation and improves serving throughout. However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory. As a consequence, one needs to rewrite the attention kernels to support paging, and implement a memory manager in the serving framework. This results in both performance and programming overheads, as well as portability challenges in adopting state-of-the-art attention kernels.
In this paper, we propose vAttention, a new approach for dynamic KV-cache memory management. In contrast to PagedAttention, vAttention stores KV-cache in contiguous virtual memory and leverages OS support for on-demand allocation of physical memory. vAttention thus enables one to use state-of-the art attention kernels out-of-the-box by adding support for dynamic allocation of physical memory without having to re-write their code. We implement vAttention in the vLLM serving stack to show that it also helps improve decode throughput by up to 1.99×1.$9$ over vLLM, and the end-to-end serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
@@ 1. Introduction
Large Language Models (LLMs) are deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (openai2022gpt4techreport, ; chowdhery2022Palm, ; bingai, ; bard, ; githubcopilot, ; amazoncodewhisperer, ; replitghostwriter, ). Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (sarathi2023, ; splitfuse2024, ; patel2023splitwise, ; distserve2024, ; tetriinfer, ; vllmsosp, ; orca, ). LLM inference happens in two phases: a compute-bound prefill phase, followed by several iterations of memory-bound decode phase. Batching is a powerful technique to boost LLM serving throughput (orca, ; vllmsosp, ; patel2023splitwise, ; sarathiserve2024, ), as it amortizes the cost of fetching model weights from HBM to the on-device SRAM during decodes.
@@
vLLM: Pioneered PagedAttention. Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention (Table 6). Further, changing block size changes the execute latency of the kernel by as much as 3×3$\times$3 × (Figure 3).
FlashAttention: PagedAttention based prefill kernel is up to $28\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower. Initial attempts to add paging support failed unit tests (fa-paged-crash, ).
TensorRT-LLM: Serving throughput dropped by more than $10\%$ in a Python front-end (tensorrt-perf-decay, ). Recommends using the C++ front-end. Even with C++, we observe up to $5\%$ higher latency in some cases with PagedAttention.
FlashInfer: PagedAttention based prefill kernel is up to $24\%$ slower than the non-paged kernel (Figure 2).
@@
Careful allocation of GPU memory is key to enabling larger batch sizes. For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for the subsequent token generation. This is called the KV-cache (orca, ; sarathi2023, ; patel2023splitwise, ) which accounts for a majority of GPU memory usage during inference. Efficiently allocating GPU memory for the KV-cache of a batch of requests is challenging for two reasons. First, the per-request KV-cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV-cache size) is not known ahead of time.
@@
Prior systems like Orca (orca, ) and FasterTransformer (fastertransformer, ) allocate GPU memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-34b-200k-hf, )). However, in practice, the number of decode tokens generated are far less; the average decode length for the chat-based sharegpt dataset is 415 tokens (sarathiserve2024, ). Therefore, static memory allocation results in severe internal fragmentation, limiting batch size and serving throughput.
@@
Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (vllmsosp, ) to mitigate KV-cache related memory fragmentation. Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens. This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trtllmgithub, ), HuggingFace TGI (hftgi, ), LightLLM (lightllm:github, ) etc.
@@
However, the PagedAttention approach faces a fundamental consequence of dynamic memory allocation: dynamically allocated memory blocks are not guaranteed to be contiguous. Note that user-level objects are allocated in virtual memory. Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the layout of KV-cache from contiguous to non-contiguous virtual memory. We argue that this approach has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
1. Requires re-writing the attention kernel (GPU code). The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient. Many implementations of deep learning operators also expect virtual contiguity of input tensors. However, by storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache. The need to re-write code is a major barrier to using new attention optimizations in production settings.
2. Adds software complexity and redundancy (CPU code). To stitch together dynamically allocated virtual memory blocks, PagedAttention forces developers to implement a memory manager in the serving framework. Typically, this requires a Block-Table manager in user space which is essentially a re-implementation of demand paging (an OS functionality) in user code.
3. Introduces performance overhead. PagedAttention adds runtime overhead in the critical path of execution in two ways. First, GPU kernels that compute attention over a non-contiguous KV-cache need to execute additional instructions. In many cases, this slows down an attention kernel by more than 20%, compared to when the KV-cache is virtually contiguous. Second, the user space memory manager can add CPU overhead contributing up to another 10% cost (§3.3).
In this paper, we instead advocate for leveraging the OS support of virtual memory and demand paging for dynamic KV-cache memory management. This approach alleviates the aforementioned issues that appear with PagedAttention.
To support our claims, we present the design and implementation of vAttention. Leveraging OS support, vAttention stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time, thereby enabling the use of state-of-the art high-performant attention kernels out-of-the box. We achieve this by using CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5).
@@
Leveraging OS support for memory allocation in an LLM serving system poses two key efficiency challenges (§5.4). First, the minimum physical memory allocation granularity supported by CUDA is 2MB. This can result in significant wasted capacity and fragmentation. We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for finer-grained physical memory allocations in multiples of 64KB. Our evaluation shows that using 64KB pages as the granularity of virtual-to-physical address translation does not affect the execution latency of attention kernels i.e., we do not find any evidence of TLB thrashing. Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the OS kernel. We introduce several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation. These optimizations hide the latency cost of memory allocation from end-users, making vAttention an efficient KV-cache memory manager.
Overall, we make the following contributions in this paper:
@@
  • We present vAttention – a memory management approach that retains the virtual contiguity of KV-cache while enabling dynamic physical memory allocation.
  • We implement vAttention in vLLM serving stack to show that it seamlessly adds dynamic memory management support to unmodified attention kernels of FlashAttention (fagithub, ) and FlashInfer (figithub, ).
@@
  • We compare vAttention against PagedAttention based alternatives of vLLM, FlashAttention and FlashInfer by evaluating Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs. Using FlashAttention’s non-paged kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput, while improving the end-to-end LLM serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
@@ 2.1. Large Language Models
Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (vllmsosp, ). Each inference request begins with a prefill phase that processes all its prompt tokens in parallel. The prefill phase produces the first output token of a request. Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (sarathi2023, ).
LLMs are built atop one of variants of the transformer architecture (attentionpaper, ). A transformer block contains two types of operators: position-wise and sequence-wise. The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator. In this paper, we primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
For the attention operator, the model first computes the query, key and value vectors from a given sequence of tokens $(x_{1},x_{2},....,x_{K})\in\mathbb{R}^{K\times E}$ where E represents the embedding size of the model. For each $x_{i}$, query, key and value vectors are computed as follows:
@@
The resulting $k_{i}$ and $v_{i}$ are appended to the key and value vectors of the prior tokens of the corresponding request, producing two matrices $K,V\in\mathbb{R}^{L^{\prime}\times(H\times D)}$ where $L^{\prime}$ represents the context length of the request seen so far, H is the number of KV heads and D is the dimension of each KV head. Then, attention is computed as follows:
@@
The attention score is computed separately for each request in the batch. Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention. Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache. A request is executed until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
Structure of the KV-cache and terminology: An LLM consists of multiple layers of the transformer block and each layer maintains its own cache of keys and values. In this paper, we refer to the cache of all transformer blocks collectively as KV-cache while using the term K-cache or V-cache for keys and values, respectively. In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request. We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a non-paged kernel.
@@ 2.2. Fragmentation and PagedAttention
To improve serving throughput, production systems rely on batching which requires careful allocation of GPU memory. This is challenging because the total context length of a request is not known in advance. Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K). vLLM shows that this strategy is prone to severe internal fragmentation. In fact, vLLM showed that prior reservation is suboptimal even if the context lengths are known in advance. This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV-cache. PagedAttention splits KV-cache into fixed-sized blocks and allocates memory for one block at a time. This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time.

### v3

@@ Abstract.
PagedAttention is a popular approach for dynamic memory allocation in LLM serving systems. It enables on-demand allocation of GPU memory to mitigate KV cache fragmentation — a phenomenon that crippled the batch size (and consequently throughput) in prior systems. However, in trying to allocate physical memory at runtime, PagedAttention ends up changing the virtual memory layout of the KV cache from contiguous to non-contiguous. Such a design leads to non-trivial programming and performance overheads.
We present vAttention — an approach that mitigates fragmentation in physical memory while retaining the virtual memory contiguity of the KV cache. We achieve this by decoupling the allocation of virtual and physical memory using CUDA virtual memory management APIs. We also introduce various LLM-specific optimizations to address the limitations of CUDA virtual memory support. Overall, vAttention is a simpler, portable, and performant alternative to PagedAttention: it supports various attention kernels out-of-the-box and improves LLM serving throughput by up to 1.23×1.23$\times$1.23 × compared to the use of PagedAttention-based kernels of FlashAttention-2 and FlashInfer.
Large language models; KV cache; fragmentation; memory management
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3669940.3707256isbn: 979-8-4007-0698-1/25/03ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Virtual memoryccs: Software and its engineering Software design techniquesccs: General and reference Performanceccs: Computing methodologies Neural networks
@@ 1. Introduction
vLLM (Kwon et al., 2023): Pioneered PagedAttention. Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention-2 (Table 7). Furthermore, changing block size changes the execution time of the kernel by as much as 1.9×1.9$\times$1.9 × (Figure 3).
FlashAttention-2 (Dao, 2023): PagedAttention-based prefill kernel is up to $37\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower. Initial attempts to add paging support failed unit tests (fa-, 2023).
FlashAttention-3 (Shah et al., 2024) / SDPA in cuDNN-9 (Nicely and NVIDIA, 2024): State-of-the-art attention kernels for NVIDIA Hopper architecture did not support PagedAttention when released.
TensorRT-LLM (trt, 2023): Serving throughput dropped by more than $10\%$ in a Python front-end (ten, 2023). Recommends using the C++ front-end. Even with C++, we observe up to $5\%$ higher latency in some cases with PagedAttention.
FlashInfer (Ye et al., 2024): PagedAttention-based prefill kernel is up to $42\%$ slower than the non-paged kernel (Figure 2).
@@
Large Language Models (LLMs) are being deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (OpenAI, 2023; Chowdhery et al., 2022; bin, 2024; bar, 2024; git, 2024; ama, 2024; rep, 2024). Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2024; Zhong et al., 2024; Hu et al., 2024; Kwon et al., 2023; Yu et al., 2022; Kamath et al., 2024).
@@
Batching is a powerful technique to boost LLM serving throughput (Yu et al., 2022; Kwon et al., 2023; Patel et al., 2024; Agrawal et al., 2024b). However, achieving a large batch size requires careful allocation of GPU memory. For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for generating subsequent tokens. This is called the KV cache (Yu et al., 2022; Agrawal et al., 2023; Patel et al., 2024) which accounts for a majority of GPU memory usage during inference. Efficiently allocating GPU memory for the KV cache is challenging for two reasons. First, the per-request KV cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV cache size) is not known ahead of time.
@@
Prior systems like Orca (Yu et al., 2022) and FasterTransformer (fas, 2024b) allocate memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-, 2024a)). However, the number of decode tokens generated are far less in practice, e.g., the average decode length for the chat-based sharegpt dataset is 415 tokens (Agrawal et al., 2024b). Therefore, static memory allocation could create severe internal fragmentation, limiting batch size and serving throughput.
@@
Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (Kwon et al., 2023) that allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens. This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems, e.g., TensorRT-LLM, HuggingFace TGI, LightLLM (trt, 2023; hft, 2024; lig, 2023) etc.
@@
However, we show that PagedAttention faces a fundamental consequence of dynamic memory allocation: dynamically allocated objects are not guaranteed to be contiguous. Note that user-level objects are allocated in virtual memory. Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the virtual memory layout of KV cache from contiguous to non-contiguous. We argue that this approach has several pitfalls (§3). First, it requires rewriting attention kernels, i.e., to enable de-referencing all tokens of the non-contiguous KV cache. Second, it forces developers to implement a memory manager in the serving framework, i.e., to stitch together dynamically allocated virtual memory blocks. Third, it adds runtime overhead in the critical path of both CPU and GPU execution. Table 1 provides empirical evidence and real-world experiences to support these arguments.
The fundamental issue with PagedAttention and prior systems is that they rely on the reservation-based memory allocation method exposed by the GPU runtime. In this method (used by cudaMalloc), the runtime allocates both virtual and physical memory on the GPU meaning that physical memory is allocated even if the corresponding virtual memory is not accessed. This is in stark contrast to OS-based demand paging (Kwon et al., 2016; Panwar et al., 2019). We show that separating the allocation of virtual and physical memory allows for more effective KV cache memory management. To support our claim, we introduce vAttention (§5) — an approach that stores KV cache in contiguous virtual memory without committing physical memory ahead-of-time. vAttention decouples the allocation of virtual and physical memory using the CUDA virtual memory management (VMM) APIs (cud, 2024).
@@
In building vAttention, we find that using CUDA VMM support for KV cache management poses two key efficiency challenges for an LLM serving system (§6). First, memory allocation using CUDA VMM APIs incurs high latency because each allocation involves a round-trip to the OS kernel. We tackle latency issues with several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation. Second, CUDA supports memory allocation only at the granularity of large pages, i.e., in multiples of 2MB. Use of large pages can create significant fragmentation. We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for smaller 64KB pages. Our evaluation shows that use of 64KB pages has no negative impact on the performance of attention kernels, i.e., we do not find any evidence of TLB thrashing. Together, these optimizations mitigate fragmentation while hiding the latency cost of on demand memory allocation, making vAttention a simpler, portable and performant alternative to PagedAttention.
Overall, we make the following contributions:
@@
  • We present vAttention – a memory management approach that retains the virtual contiguity of KV cache while enabling dynamic allocation of physical memory. Our implementation of vAttention in vLLM seamlessly adds dynamic memory allocation support to various unmodified attention kernels.
@@
  • We compare vAttention against PagedAttention-based alternatives of vLLM, FlashAttention-2 and FlashInfer on Yi-6B, Llama-3-8B and Yi-34B with 1-2 A100 GPUs. Using FlashAttention-2’s non-paged attention kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput. In long-context scenarios, it also improves the end-to-end LLM serving throughput by up to 1.18×1.18$\times$1.18 × and 1.23×1.23$\times$1.23 × over PagedAttention based kernels of FlashAttention-2 and FlashInfer
  • We demonstrate the portability benefit of vAttention with the recently launched FlashAttention-3 kernel (FA3 (Shah et al., 2024)). FA3 is optimized for the NVIDIA Hopper architecture and was not released with PagedAttention support. vAttention supports FA3 out-of-the-box, leading to 1.261.5×1.26-1.5$\times$1.26 - 1.5 × higher throughout over PagedAttention based FlashAttention-2.
@@ 2.1. Large Language Models
Given an input sequence, an LLM predicts the probability of an output sequence: a sequence is a set of tokens (Kwon et al., 2023). Each inference request begins with a prefill phase that processes prompt tokens in parallel and produces the first output token of the request. Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (Agrawal et al., 2023).
LLMs are built atop one of the variants of the transformer architecture (Vaswani et al., 2017); an LLM consists of multiple transformer blocks. Internally, a transformer block contains two types of operators: position-wise and sequence-wise (Yu et al., 2022). The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator. We primarily focus on attention since it is the primary consumer of GPU memory in LLM inference. Table 2 summarizes the notations used in the paper.
For the attention operator, the model computes the query, key and value vectors from a given sequence of tokens $(x_{1},x_{2},....,x_{K})\in\mathbb{R}^{K\times E}$ where E represents the embedding size of the model. For each $x_{i}$, query, key and value vectors are computed as follows:
@@
The resulting $k_{i}$ and $v_{i}$ are appended to the key and value vectors of the prior tokens of the corresponding request, producing two matrices $K,V\in\mathbb{R}^{L^{\prime}\times(H\times D)}$ where $L^{\prime}$ represents the context length of the request seen so far, H is the number of KV heads of the model on a worker and D is the dimension of each KV head. Then, attention is computed as follows:
@@
Symbol Definition
$N$ Number of layers on a particular worker
$H$ Number of KV heads on a particular worker
$D$ Dimension of each attention head
$P$ Number of bytes needed to store one element
$B$ Maximum batch size
$L$ Maximum context length supported by the model
$L^{\prime}$ Context length of a request seen thus far
Table 2. Notations used in the paper.
The attention score is computed separately for each request in the batch. A request executes until the model generates a special end-of-sequence token or reaches the maximum context length for the request. Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention. Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations. We refer to this state of all layers collectively as KV cache. In systems prior to PagedAttention, the K cache (or V cache) at each layer of a worker is typically allocated as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request.
@@ 2.2. Fragmentation and PagedAttention
Serving LLMs with high throughput requires careful allocation of GPU memory. This is challenging because the total context length of a request is not known in advance. Serving systems worked around this challenge by pre-reserving KV cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K). vLLM shows that this strategy is prone to severe internal fragmentation. In fact, vLLM showed that prior reservation is suboptimal even if the context lengths are known in advance. This is because the per-request KV cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV cache. PagedAttention splits KV cache into fixed-sized blocks and allocates memory for one block at a time. This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time.
GPUs and page sizes: NVIDIA GPUs support multiple page sizes in the hardware (Pratheek et al., 2022; Nayak et al., 2021; Zhang et al., 2023a; pas, 2024). A single call to a CUDA VMM API can allocate one or more physical pages, which we refer to as a page-group. We use page-groups to support multiple allocation granularities for the KV cache, similar to how multiple page sizes are commonly used in conventional OS-based virtual memory systems (Kwon et al., 2016; Panwar et al., 2019).
3. Issues with the PagedAttention Approach
Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention implements demand paging in user space whereas conventional demand paging is transparent to applications. This section elaborates on issues that arise with such an approach.
3.1. Requires Re-writing the Attention Kernel
Conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory. By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV cache blocks. Writing correct and performant GPU kernels can be challenging for most programmers (fa-, 2023).
Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (Dao et al., 2024; Dao, 2023; fla, 2023; Hong et al., 2024; Ye et al., 2024; Zhang et al., 2023b; Bikshandi and Shah, 2023; Ainslie et al., 2023; Shazeer, 2019; Child et al., 2019; Beltagy et al., 2020; Kamath et al., 2024), and this trend is likely to continue. In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation. Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage. To provide an example, Table 7 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention-2 kernel (Ainslie et al., 2023).